OpenStudy (anonymous):

Please help.. Solve for t: 3(2^t)=11 Solve for x: e^2x= 8 Solve for x: ln(5x-1)= 6

OpenStudy (stapeleliz):

To learn how to solve log equations, try here: Solving Logarithmic Equations http://www.purplemath.com/modules/solvelog.htm To learn how to solve exponential equations, try here: Solving Exponential Equations http://www.purplemath.com/modules/solvexpo.htm

OpenStudy (turingtest):

3(2^t)=11 2^t=11/3 t=log_2(11/3) (that's log base 2) e^(2x)= 8 2x=ln(8) x=(1/2)ln(8) ln(5x-1)= 6 5x-1=e^6 5x=1+e^6 x=(1/5)(1+e^6)